Anela4ever:

Select all the equations that are NOT equivalent to p – 7 = 22. p – p – 7 = 22 – p p – 7 – p = 22 – p p – 7 + 22 = 22 – 7 p – 7 + 7 = 22 – 22 p – 7 + 7 = 22 + 7

Glorious:

It's all of them. None of the equations are equivalent to p - 7 = 22.

Anela4ever:

I can only pick two out of all of them

fwval:

@anela4ever wrote:
I can only pick two out of all of them
the first 2 are not equivalent to p - 7 = 22

Anela4ever:

so then would it be the last two

fwval:

p – p – 7 = 22 – p p – 7 – p = 22 – p

fwval:

these both ^

Anela4ever:

thanks for the help but one of them was wrong